# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ 4. Rigidity and existence of the discrete conformal structure (A2)
By Remark 1.7, we can set $C\equiv 0$ in the discrete conformal structure (A2).By Remark 1.7, we can set $C\equiv 0$ in the discrete conformal structure (A2). In this case, the discrete conformal structure (A2) is the $(-1,-1,0)$ type generalized circle packings in Guo-Luo [5]. Its rigidity and existence have been proved by Guo-Luo [5], which is as follows.
@@
([5]) Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary and the weight $\eta\in\mathbb{R}_{>0}^{E}$, the discrete conformal structure $f\in\Omega(\eta)$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$. Furthermore, the image of $K$ is $\mathbb{R}_{>0}^{N}$.
In this section, we reprove the rigidity part of Theorem 4.1. In [5], Guo-Luo spent lots of space to prove the symmetry and negative definiteness of Jacobian. By (2) and (3), we simplify Guo-Luo’s proof, which is similar to that of Theorem 3.2.
@@
The admissible space of the discrete conformal factor $f$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ given by (7) is defined to be
$\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}|\cosh l_{rs}=-\cosh(f_{r}-f_{s})+\eta_{rs}e^{f_{r}+f_{s}}>1,\ \forall\{r,s\}\subseteq\{i,j,k\}\}.$
@@
By Theorem 1.6, we have $d>0$ for the discrete conformal structure (A2). By a change of variables in Remark 2.23, i.e., $\frac{df_{r}}{du_{r}}=e^{f_{r}}$ and $u_{r}=-e^{-f_{r}}$. Thus
@@
Since $u_{r},u_{s}\in\mathbb{R}_{<0}$, then $\eta_{rs}>0$ and $u_{r}+u_{s}>-\sqrt{2\eta_{rs}}$. Therefore, the admissible space $\Omega_{ijk}(\eta)$ of $f$ is transferred to the following admissible space $\mathcal{U}_{ijk}(\eta)$ of $u$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$,
$\mathcal{U}_{ijk}(\eta)=u(\Omega_{ijk}(\eta))=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{<0}^{3}|u_{r}+u_{s}>-\sqrt{2\eta_{rs}},\ \forall\{r,s\}\subseteq\{i,j,k\}\},$
which is a convex polytope. Then the admissible space of $u$ on a weighted triangulated surface $(\Sigma,\mathcal{T},\eta)$ can be defined as
$\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta),$
which is also a convex polytope.
@@
Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary and the weight $\eta\in\mathbb{R}_{>0}^{E}$. For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\mathcal{U}_{ijk}(\eta)$, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is non-degenerate, symmetric and negative definite. As a result, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
@@
For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r}$ and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$. Since $\frac{df_{r}}{du_{r}}=e^{f_{r}}$, then
$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}e^{f_{i}}&0&0\\ 0&e^{f_{j}}&0\\ 0&0&e^{f_{k}}\\ \end{array}\right):=Q_{4}.$
Hence
where $Q_{1}$ is defined by (36) and $Q_{2}$ is defined by (38). Moreover, $\det Q_{1}>0$.
Since $d>0$ for the discrete conformal structure (A2) in Theorem 1.6, then $\det Q_{2}>0$ by (39). Hence,
$\det\Lambda_{ijk}=-\frac{1}{A^{3}}\sinh l_{i}\sinh l_{j}\sinh l_{k}\cdot\det Q_{1}\cdot\det Q_{2}\cdot\det Q_{4}<0,$
which implies $\Lambda_{ijk}$ is non-degenerate. The symmetry of $\Lambda_{ijk}$ follows from Lemma 2.22. Set $a_{rrs}=e^{f_{r}}\cdot\coth d_{rs}>0$. Define
The rest of the proof is similar to that of Theorem 3.2, so we omit it. Q.E.D.
The following theorem gives the rigidity of discrete conformal structure (A2), which is the case $\mathrm{(ii)}$ in Theorem 1.9. As the proof is similar to that of Theorem 3.4, we omit it here.
@@
Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary and the weight $\eta\in\mathbb{R}_{>0}^{E}$. For the discrete conformal structure (A2), the function $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{E}_{>0}$. In particular, the map from $f$ to $K$ is a smooth embedding.
5. Rigidity and existence of the mixed discrete conformal structure I
In this section, we consider the rigidity and existence for the mixed discrete conformal structure $\mathrm{I}$, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}\in F$ are induced by (7) and (10). By Remark 1.7, we can set $C\equiv 0$ for (7) and (10). As mentioned in Remark 1.8, if $d_{rs}<0$, then $l_{rs},l_{rt}$ are given by (10) and $l_{st}$ is given by (7), where $\{r,s,t\}=\{i,j,k\}$. Without loss of generality, we can assume $d_{ij}<0$ and $d_{ik}<0$. Then for a right-angled hyperbolic hexagon $\{ijk\}\in F$, the edge lengths $l_{ij},l_{ik}$ are given by (10) and the edge length $l_{jk}$ is given by (7), i.e.,
(43) $\cosh l_{ij}$ $=\cosh(f_{j}-f_{i})+\zeta_{ij}e^{f_{i}+f_{j}},$
$\cosh l_{ik}$ $=\cosh(f_{k}-f_{i})+\zeta_{ik}e^{f_{i}+f_{k}},$
@@ 5. Rigidity and existence of the mixed discrete conformal structure I
Here $\eta_{ij}\in\mathbb{R}_{>0}$ and $\zeta_{ij},\zeta_{ik}\in\mathbb{R}$.
5.1. Rigidity of the mixed discrete conformal structure $\mathrm{I}$
The admissible space of the discrete conformal factor $f$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43) is defined to be
$\Omega_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}|l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
Combining (30) and Remark 2.23 gives $u_{i}=e^{-f_{i}}\in\mathbb{R}_{>0}$, $u_{j}=-e^{-f_{j}}\in\mathbb{R}_{<0}$ and $u_{k}=-e^{-f_{k}}\in\mathbb{R}_{<0}$. Thus

### v3

@@ 4. Rigidity and existence of the discrete conformal structure (A2)
By Remark 1.6, we set $\alpha\equiv-1$. Then (6) can be written as
(43) $\cosh l_{ij}=\sqrt{(e^{2f_{i}}-1)(e^{2f_{j}}-1)}+\eta_{ij}e^{f_{i}+f_{j}}.$
For a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ given by (43), the admissible space of the discrete conformal factors $f$ is defined as
(44) $\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}_{>0}^{3}\mid l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
Since $\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{e^{2f_{i}}-1}$, it follows from Remark 2.22 (iv) that $\tan u_{i}=\sqrt{e^{2f_{i}}-1}$ for $u_{i}\in(0,\frac{\pi}{2})$ and $-\frac{1}{\tan u_{i}}=\sqrt{e^{2f_{i}}-1}$ for $u_{i}\in(-\frac{\pi}{2},0)$. For simplicity, we consider $u\in(-\frac{\pi}{2},0)^{N}$. Consequently,
(45) $\cosh l_{ij}=\frac{1}{\tan u_{i}\tan u_{j}}+\eta_{ij}\frac{1}{\sin u_{i}\sin u_{j}}>1,$
which implies $\cos(u_{i}+u_{j})>-\eta_{ij}$. This further yields $u_{i}+u_{j}\in(-\pi,0)$ when $\eta_{ij}\in[1,+\infty)$, and $u_{i}+u_{j}>-\arccos\eta_{ij}$ when $\eta_{ij}\in(-1,1)$. Hence, the admissible space $\Omega_{ijk}(\eta)$ is transformed into the following admissible space $\mathcal{U}_{ijk}(\eta)$ of $u$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$:
(46) $\mathcal{U}_{ijk}(\eta)=u(\Omega_{ijk}(\eta))=\{(u_{i},u_{j},u_{k})\in(-\frac{\pi}{2},0)^{3}\mid\cos(u_{r}+u_{s})>-\eta_{rs},\,\forall\{r,s\}\subseteq\{i,j,k\}\},$
where $\eta_{rs}\in(-1,+\infty)$. Hence, $\mathcal{U}_{ijk}(\eta)$ is a convex polytope. As a result, the admissible space $\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$ is also a convex polytope on $(\Sigma,\mathcal{T},\eta)$.
@@
Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary, where $\eta\in(-1,+\infty)^{E}$. For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\mathcal{U}_{ijk}(\eta)$, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is symmetric and negative definite. Consequently, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
Proof.
For simplicity, we adopt the notations introduced in Subsection 3.2. Define
$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}\sqrt{e^{2f_{i}}-1}&0&0\\ 0&\sqrt{e^{2f_{j}}-1}&0\\ 0&0&\sqrt{e^{2f_{k}}-1}\\ \end{array}\right):=Q_{4}.$
Then $\Lambda_{ijk}$ can be expressed as
where $Q_{1}$ and $Q_{2}$ are defined by (36) and (38), respectively. Furthermore, $\det Q_{1}>0$. Since $d>0$ for the discrete conformal structure (A2) in Theorem 1.5, it follows from (39) that $\det Q_{2}>0$. Consequently, $\det\Lambda_{ijk}<0$, which implies $\Lambda_{ijk}$ is non-degenerate. The symmetry of $\Lambda_{ijk}$ follows from Lemma 2.21. Set $a_{rrs}=\sqrt{e^{2f_{r}}-1}\cdot\coth d_{rs}>0$. Define
The remaining part of the proof is analogous to that of Theorem 3.2, and thus is omitted here.
Q.E.D.
The following theorem establishes the rigidity of the discrete conformal structure (A2), which corresponds to the rigidity part of Theorem 1.8 (ii). As its proof is analogous to that of Theorem 3.4, it is omitted here.
Theorem 4.2.
@@
Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary, where $\eta\in(-1,+\infty)^{E}$. The discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
The following theorem characterizes the image of $K$ for the discrete conformal structure (A2), which corresponds to the existence part of Theorem 1.8 (ii).
Theorem 4.3.
Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary, where $\eta\in(-1,0]^{E}$. The image of $K$ is >$\mathbb{R}^{N}_{>0}$.
To prove Theorem 4.3, we need the following lemma. The proof is nearly identical to that of Lemma 3.6, and is thus omitted here.
Lemma 4.4.
Under the same assumptions as those in Theorem 4.2, for a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\Omega_{ijk}(\eta)$, if one of the following conditions is satisfied
(i):
$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$,
(ii):
$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=c$,
(iii):
$\lim f_{i}=+\infty,\ \lim f_{j}=a,\ \lim f_{k}=b$,
where $a,b,c$ are positive constants. Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$.
Proof of Theorem 4.3: Let $X$ be the image of $K$. The openness of $X$ in >$\mathbb{R}^{N}_{>0}$ is a consequence of the smooth injectivity of the map, a property that plays a pivotal role in establishing the rigidity theorem. To prove the closedness of $X$ in >$\mathbb{R}^{N}_{>0}$, we show that for any sequence $\{f^{(m)}\}\in\Omega(\eta)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\eta)$ satisfying $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$, there exists a subsequence, still denoted by $\{f^{(m)}\}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}\in\Omega(\eta)$.
To derive a contradiction, suppose there exists a subsequence $\{f^{(m)}\}$ such that $\lim_{m\rightarrow+\infty}f^{(m)}$ lies on the boundary of $\Omega(\eta)$. By (44), the boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ consists of the following three parts:
(i):
There exists an ideal edge $\{rs\}\in E$ such that $l_{rs}=0$;
(ii):
There exists $i\in B$ such that $f_{i}^{(m)}\rightarrow+\infty$;
(iii):
There exists $i\in B$ such that $f_{i}^{(m)}\rightarrow 0^{+}$.
The proof for the case (i) is identical to that of Theorem 3.5 and is thus omitted here. For the case (ii), by Lemma 4.4, each generalized angle incident to the boundary component $i\in B$ converges to 0. Consequently, $\lim_{m\rightarrow+\infty}K_{i}^{(m)}=0$. This contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$. For the case (iii), if $f_{i}^{(m)}\rightarrow 0^{+}$, then $\cosh l_{ij}=\sqrt{(e^{2f_{i}}-1)(e^{2f_{j}}-1)}+\eta_{ij}e^{f_{i}+f_{j}}\rightarrow\eta_{ij}e^{f_{j}}\geq 1$. However, since $\eta_{ij}\in(-1,0]$, it follows that $\eta_{ij}e^{f_{j}}\leq 0$, which is a contradiction. Q.E.D.
Remark 4.5.
If $\eta=-\cos\Phi\in(-1,1]^{E}$ and $e^{f}=\cosh r$, then the discrete conformal structure (A2) reduces to the $(-1,-1,1)$-type generalized circle packings introduced by Guo-Luo in [5], where the rigidity of such generalized circle packings was also established. Consequently, Theorem 4.2 generalizes the rigidity result for Guo-Luo’s $(-1,-1,1)$-type generalized circle packings in [5]. Guo-Luo further asserted that the existence holds for $\eta\in(-1,1]^{E}$. However, their proof in [5] contains a gap when $\eta\in(0,1)^{E}$, because they did not account for the boundary behavior corresponding to $r\rightarrow 0^{+}$. As a result, Theorem 4.3 holds for $\eta\in(-1,0]^{E}$ rather than $\eta\in(-1,1]^{E}$.
5. Rigidity and existence of the discrete conformal structure (A3)
By Remark 1.6, we set $C\equiv 0$ in the discrete conformal structure (A3). In this case, the discrete conformal structure (A3) coincides precisely with the $(-1,-1,0)$-type generalized circle packings introduced by Guo-Luo in [5]. Guo-Luo [5] have established the rigidity and existence results for such structures, which can be stated as follows.
Theorem 5.1.
([5]) Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary, where $\eta\in\mathbb{R}_{>0}^{E}$. The discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$. Furthermore, the image of $K$ is $\mathbb{R}_{>0}^{N}$.
In this section, we present a concise proof for the rigidity part of Theorem 5.1, which simplifies the argument given by Guo-Luo in [5]. For the proof of the existence part, we refer the reader to [5].
@@ 5. Rigidity and existence of the discrete conformal structure (A3)
For a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ given by (7), the admissible space of the discrete conformal factor $f$ is defined as
$\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}\mid\cosh l_{rs}=-\cosh\,(f_{r}-f_{s})+\eta_{rs}e^{f_{r}+f_{s}}>1,\,\forall\{r,s\}\subseteq\{i,j,k\}\}.$
@@
By Theorem 1.5, for the discrete conformal structure (A3), we have $d>0$. By the variable transformation in Remark 2.22, specifically, $\frac{df_{r}}{du_{r}}=e^{f_{r}}$ and $u_{r}=-e^{-f_{r}}$, we derive
@@
Since $u_{r}\in\mathbb{R}_{<0}$ and $u_{s}\in\mathbb{R}_{<0}$, it follows that $\eta_{rs}>0$ and $u_{r}+u_{s}>-\sqrt{2\eta_{rs}}$. Consequently, for a right-angled hyperbolic hexagon $\{ijk\}\in F$, the admissible space $\Omega_{ijk}(\eta)$ of $f$ is transformed into the following admissible space $\mathcal{U}_{ijk}(\eta)$ of $u$:
(47) $\mathcal{U}_{ijk}(\eta)=u(\Omega_{ijk}(\eta))=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{<0}^{3}\mid u_{r}+u_{s}>-\sqrt{2\eta_{rs}},\,\forall\{r,s\}\subseteq\{i,j,k\}\},$
which is a convex polytope. Consequently, the admissible space $\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$ is also a convex polytope on $(\Sigma,\mathcal{T},\eta)$.
@@
Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary, where $\eta\in\mathbb{R}_{>0}^{E}$. For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\mathcal{U}_{ijk}(\eta)$, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is symmetric and negative definite. Consequently, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
@@
For simplicity, we adopt the notations introduced in Subsection 3.2. Since $\frac{df_{r}}{du_{r}}=e^{f_{r}}$, it follows that
$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}e^{f_{i}}&0&0\\ 0&e^{f_{j}}&0\\ 0&0&e^{f_{k}}\\ \end{array}\right):=Q_{5}.$
Then $\Lambda_{ijk}$ can be expressed as
where $Q_{1}$ and $Q_{2}$ are defined by (36) and (38), respectively. Furthermore, $\det Q_{1}>0$. Since $d>0$ for the discrete conformal structure (A3) in Theorem 1.5, it follows from (39) that $\det Q_{2}>0$. Consequently, $\det\Lambda_{ijk}<0$, which implies $\Lambda_{ijk}$ is non-degenerate. The symmetry of $\Lambda_{ijk}$ follows from Lemma 2.21. Set $a_{rrs}=e^{f_{r}}\cdot\coth d_{rs}>0$. Define
$Q:=Q_{1}Q_{2}Q_{5}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right)\\ \times\left(\begin{array}[]{ccc}0&a_{jjk}&a_{kkj}\\ a_{iik}&0&a_{kki}\\ a_{iij}&a_{jji}&0\\ \end{array}\right).$
The remaining part of the proof is analogous to that of Theorem 3.2, and thus is omitted here.
Q.E.D.
The following theorem establishes the rigidity of the discrete conformal structure (A3), which corresponds to the rigidity part of Theorem 1.8 (iii). As its proof is similar to that of Theorem 3.4, it is omitted here.
@@
Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary, where $\eta\in\mathbb{R}_{>0}^{E}$. The discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
6. Rigidity and existence of the mixed discrete conformal structure III
In this section, we investigate the rigidity and existence of the mixed discrete conformal structure III, specifically, an ideally triangulated surface with boundary admitting both discrete conformal structures (A3) and (B3). All right-angled hyperbolic hexagons are either Type-III or a combination of Type-III right-angled hyperbolic hexagons and those with edge lengths given by (7). Here, a right-angled hyperbolic hexagon is Type-III if its edge lengths are given by both (7) and (10). By Remark 1.6, we set $C\equiv 0$ for both (7) and (10).
As noted in Remark 1.7, if $d_{rs}<0$, then $d_{rt}<0,\ d_{sr}>0,\ d_{tr}>0,\ d_{st}>0,\ d_{ts}>0$, and $l_{rs},l_{rt}$ are given by (10) and $l_{st}$ is given by (7), where $\{r,s,t\}=\{i,j,k\}$. Without loss of generality, we assume $d_{ij}<0$ in this section. This implies $d_{ik}<0$, $d_{ji}>0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$. For the Type-III right-angled hyperbolic hexagon $\{ijk\}\in F$, the edge lengths $l_{ij},l_{ik}$ are given by (10), and $l_{jk}$ is given by (7), namely,
(48) $\cosh l_{ij}$ $=\cosh(f_{j}-f_{i})+\eta_{ij}e^{f_{i}+f_{j}},$
$\cosh l_{ik}$ $=\cosh(f_{k}-f_{i})+\eta_{ik}e^{f_{i}+f_{k}},$
@@ 6. Rigidity and existence of the mixed discrete conformal structure III
Here $\eta_{ij}\in\mathbb{R}$, $\eta_{ik}\in\mathbb{R}$, and $\eta_{jk}\in\mathbb{R}_{>0}$.
6.1. Rigidity of the mixed discrete conformal structure III
For a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (48), the admissible space of the discrete conformal factor $f$ is defined as
$\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}\mid l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
Combining (29) and Remark 2.22 yields $u_{i}=e^{-f_{i}}\in\mathbb{R}_{>0}$, $u_{j}=-e^{-f_{j}}\in\mathbb{R}_{<0}$, and $u_{k}=-e^{-f_{k}}\in\mathbb{R}_{<0}$. By (48), we have